OpenStudy (anonymous):

y = –x – 4 and y = 4x + 1 (–1, –5) (–1, –3) (1, –3) (1, 5)

OpenStudy (shubhamsrg):

subtract both eqns,,solve for x..and then solve for y..

OpenStudy (radar):

y=y -x-4=4x+1 -5x=5 x=-1 Now substitute for x in either equation and solve for y.